You answered all 100 questions on a test. You had 74 more correct answers than incorrect answers. How many correct answers did you have?
I know what the answer is, however, I need to know how that answer is figured showing the problems, thinking, & strategies used.
Thank you
Cherylin

Answer by ankor@dixie-net.com(22740)

You answered all 100 questions on a test.
You had 74 more correct answers than incorrect answers.
How many correct answers did you have?
:
Let c = no. of correct answers
Let w = no. of wrong answers
:
Write an equation for each statement:
:
"You answered all 100 questions on a test."
c + w = 100
:
" You had 74 more correct answers than incorrect answers."
w = c - 74
;
Replace w with (c-74) in the 1st equation
c + (c-74) = 100
c + c = 100 + 74
2c = 174
c = 174/2
c = 87 correct answers
;
:
Check this: 87 - 74 = 13 wrong answer
13 + 87 = 100 confirms our solution of c = 87